# PAC-Chernoff Bounds: Understanding Generalization in the Interpolation Regime

### v3

@@ 6.2 Data Augmentation
Data augmentation is a widely-used technique in machine learning that enhances the robustness and generalization ability of models, particularly in the field of computer vision (Shorten and Khoshgoftaar, 2019). By applying various transformations like rotations, reflections, scaling, and translations to existing data samples, data augmentation artificially expands the dataset. This process helps the learned models to become invariant or less sensitive to these transformations, leading to improved performance on unseen data.
@@
Recent research has delved deeply into the theoretical foundations of data augmentation (DA), with significant contributions coming from studies such as Chen et al. (2020) and Lyle et al. (2020). These works build on the assumption than the set of transformations $G$ forms a group and that these transformations do not alter the data-generating distribution, also assuming an equality in distribution assumption. Within this context, these studies illustrate how DA can decrease variance and improve bounds on generalization error. Lyle et al. (2020) further presents a PAC-Bayes bound based on these premises, showing that data augmentation can lower the bound’s complexity measure. However, these studies encounter a critical challenge: the equality in distribution assumption does not hold in reality, as it implies that any model ${\boldsymbol{\theta}}$ would yield identical expected losses on both transformed and non-transformed input data, a premise contradicted by evidence shown in Figure 6, where two models exhibit deteriorating performance as the data is increasingly transformed. Additionally, the PAC-Bayes bound introduced by Lyle et al. (2020) is not proven to be tight for interpolators. Thus, as highlighted in the introduction, merely demonstrating that data augmentation reduces the complexity term of a bound, which might be vacuous, does not assure a lower generalization error for an interpolator.
@@
In this section, we demonstrate that our theoretical framework offers a more nuanced explanation. First, we will illustrate that data augmentation defines a new, more concentrated empirical loss. This finding is consistent with prior studies indicating that data augmentation decreases the variance of the empirical loss (Chen et al., 2020; Lyle et al., 2020). Second, we will explore how the complexity term of the PAC-Chernoff bound, introduced in Theorem 4.1, becomes smaller with the application of data augmentation. As this bound is also perfectly tight for interpolators trained on augmented datasets, we can deduce that the generalization error of interpolators under data augmentation is reduced, in opposite to the non-tight bounds used in earlier research (Lyle et al., 2020).
@@
As argued, for example, by Chen et al. (2020), data augmentation is equivalent to optimize, using Monte-Carlo estimates, the so-called data-augmented loss, denoted as $\ell_{G}$, which is the result of averaging the standard loss over all transformations obtained from a given input ${\boldsymbol{x}}$,
@@ 7 Over-parameterization and Smooth Interpolation
Many theoretical contributions have established that over-parameterization is directly related to the strong performance of interpolators (Neyshabur et al., 2018; Bubeck et al., 2021). Specifically, the celebrated study by (Bubeck and Sellke, 2023) [^2] Outstanding Paper Award at NeurIPS’21., building on an isoperimetry assumption, demonstrates that any model capable of interpolating training data beneath the noise threshold requires a (Euclidean) Lipschitz constant of the order at least $\sqrt{nd/p}$, where $d$ represents the data’s ambient dimension. We sketch this result using our notation: We say a model ${\boldsymbol{\theta}}$ is interpolating below the noise threshold if $\hat{L}(D,{\boldsymbol{\theta}})\leq L^{\star}:=\min_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}L({\boldsymbol{\theta}})$. Similarly, we denote $Lip({\boldsymbol{\theta}})$ the Lipschitz constant of the model ${\boldsymbol{\theta}}$, as defined in Equation (6). Bubeck and Sellke (2023) (Theorem 1) can be informally stated as: under an isoperimetry assumption, for any $\epsilon\in(0,L^{\star})$ and any $\delta\in(0,1)$, with high probability $1-\delta$ over $D\sim\nu^{n}$, for all ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}$,
@@
The primary advantage of the aforementioned findings are their ability to establish a link between over-parameterization, interpolation, and our prior analyses about model’s smoothness and its rate function. This enables us to show how Bubeck and Sellke (2023)’s analysis of smooth interpolation, using isoperimetry and Lipschitz continuity, represents merely one of several approximations to the characterization presented in Theorem 7.1. As described Section 5, the norm of parameters, the distance from initialization, the input-gradient norm, and the Lipschitz constant of a model, serve as proxies of the (inverse) rate function of a model and, consequently, of its smoothness. We first proceed to show how Bubeck and Sellke (2023)’s characterization of smooth interpolation can be directly deduced from Theorem 7.1 under the same isoperimetry assumption [^3] This result can also be derived under log-concave assumptions using the same techniques used in Proposition 5.5 and Equation (6)..
@@
$\text{if}\quad\hat{L}(D,{\boldsymbol{\theta}})\leq\epsilon\quad\text{then}\quad Lip({\boldsymbol{\theta}})\geq\sqrt{\frac{nd}{2c(p\ln k-\ln\delta)}}(L^{\star}-\epsilon)\,.$
We can similarly derive a bound for interpolators with a giving norm or distance-from-initialization.
@@
$\text{if}\quad\hat{L}(D,{\boldsymbol{\theta}})\leq\epsilon\quad\text{then}\quad\|{\boldsymbol{\theta}}-{\boldsymbol{\theta}}_{0}\|_{2}\geq\sqrt{\frac{n}{8M(p\ln k-\ln\delta)}}(L^{\star}-\epsilon)\,.$
Corollary 7.2 and 7.3 imply that, as the number of samples $n$ increases, to have smooth interpolators with a small Lipschitz constant, or a small parameter norm, or a small distance from initialization, we need to increase the number of parameters of the model class too. In any case, all these results are approximations of the general result given in Theorem 7.1. Note that a similar result could be derived in the context of input-gradient norms using Proposition 5.5.
@@
7.1 Double-Descent and Smooth Interpolation
The existing literature consistently demonstrates that interpolators with a larger number of parameters tend to perform better. A finding that defies traditional beliefs from classical statistical learning theory. Traditionally, it was assumed that increasing the number of parameters in a model would lead to higher overfitting and, consequently, poor generalization performance. However, this perspective has been challenged by the phenomenon of the double-descent curve (Belkin et al., 2019), which exemplifies a shift in dynamics as models enter the interpolation regime. In this regime, an increase in parameters paradoxically leads to improved performance. This counter-intuitive behavior suggests that once models begin to interpolate, their generalization performance improves as they grow larger.
Parameters Train Acc. Test Acc. Test NLL
$4$k $43.84\%$ $42.84\%$ $1.56$
$32$k $70.94\%$ $59.76\%$ $1.17$
$85$k $86.72\%$ $62.96\%$ $1.35$
$163$k $97.72\%$ $63.78\%$ $1.82$
$266$k $99.64\%$ $65.01\%$ $2.12$
$395$k $100.0\%$ $65.85\%$ $2.34$
$548$k $100.0\%$ $66.98\%$ $2.12$
$727$k $100.0\%$ $69.71\%$ $1.81$
$931$k $100.0\%$ $69.14\%$ $1.80$
$1161$k $100.0\%$ $69.76\%$ $1.73$
Figure 9: Illustration of the double-descent phenomenon: Evolution of the train and test loss of nested convolutional neural networks with increasing number of parameters. Increasing the number of parameters promotes models with worse population loss until the critical region; further increasing the number of parameters decreases the population error. The point where models reach the interpolation regime is marked with a black line. The training and test loss of models are marked with points and their smoothed curve is shown. Only a few of the total models are shown in the table.
Open Question 8 Why does the generalization performance of interpolators improve with an increasing number of parameters?
Figure 10: Rate functions obtained by interpolating CNNs with 3k-1069k parameters on Cifar10. Parameters are increased by increasing the number of channels in the architecture. Models are divided so that the rate function decrease on left and increase on right.
The following result, viewed as an extension of our finding presented in Theorem 4.5, offers a partial explanation for the previously posed question. More precisely, it demonstrates that if the interpolators become progressively smoother (higher rate function), their generalization performance will improve, despite being part of a model class characterized by a greater number of parameters.
Theorem 7.4.
Let ${\boldsymbol{\Theta}}\subset{\boldsymbol{\Theta}}^{\prime}$ be nested model classes with $p<p^{\prime}$ parameters respectively. For any $\epsilon>0$, with h.p. $1-\delta$ over $D\sim\nu^{n}$, for any ${\boldsymbol{\theta}}^{\prime}\in{\boldsymbol{\Theta}}^{\prime}$, ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}$, simultaneously
[…]
$\Downarrow$
Figure 10 (right) illustrates the rate functions of the interpolators from Figure 9, highlighting how larger interpolators become increasingly smoother, meaning their rate functions are progressively higher. In fact, by reversing the implication of Theorem 7.4, we can infer that if a larger interpolator exhibits better generalization performance, then, it cannot be less smooth than a smaller interpolator:
Corollary 7.5.
Let ${\boldsymbol{\Theta}}\subset{\boldsymbol{\Theta}}^{\prime}$ be nested model classes with $p<p^{\prime}$ parameters respectively. For any $\epsilon>0$, with h.p. $1-\delta$ over $D\sim\nu^{n}$, for any ${\boldsymbol{\theta}}^{\prime}\in{\boldsymbol{\Theta}}^{\prime}$, ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}$, simultaneously,
$\Downarrow$
As a result, the previous open question can be rephrased as follows,
Open Question 9 Why does the smoothness of interpolators improve with an increasing number of parameters?
Based on the two previous results, addressing Open Question 7.1 is equivalent to answering Open Question 7.1. In other words, if we can elucidate why the rate function of the interpolators is increasingly higher as we increase the number of parameters, Theorem 7.4 will effectively explain why the generalization performance of interpolators enhances as the number of parameters grows, and hence, the double descent phenomena.
Our hypothesis for explaining this phenomenon posits that larger neural networks are better at capturing invariances in the data, a notion supported both theoretically (Bronstein et al., 2021) and empirically (Goodfellow et al., 2009). This capability leads to smoother models with higher rate functions, as outlined in Section 6. Consequently, Theorem 7.4 elucidates why the generalization performance of interpolators improves with an increase in parameters. The experimental setup presented in Figures 9 and 10 support this understanding.
Finally, it is also quite interesting to look at the models on the left part of the interpolation threshold of Figure 9. Surprisingly, the sequence of rate functions, displayed in Figure 10 (left), get increasingly lower as the size of the mode class increases, just the opposite of what happens in Figure 10 (right). However, explaining this behavior in the non-interpolation regime is out of the scope of this work, as the presented PAC-Chernoff bound is not tight in that regime.
@@ 8 Conclusions and Limitations
In this work, we started discussing how a growing literature (Zhang et al., 2017; Nagarajan and Kolter, 2019c; Wang et al., 2024; Gastpar et al., 2023) is clearly suggesting that bounds that solely depend on the training data are provably vacuous for over-parameterized model classes, and points to the need of exploring alternative bounds to be able to understand generalization in deep learning. We then argued in favor of distribution-dependent bounds, as bounds which directly depend on the data generating distribution, instead of the training data sample. A distribution-dependent PAC-Chernoff bound, outlined in Theorem 4.1, is introduced, which shows how this bound is perfectly tight for any interpolator. Based on the complexity measure of this bound —the inverse rate function, a distribution-dependent quantity— a novel definition of smoothness is proposed in Definition 3, which is able to provide a clear answer to the previously unanswered question of which interpolators generalize more effectively, as shown in Theorem 4.5.
In Section 5, we showed how the complexity measure of our PAC-Chernoff bound defines an regularizer that provides near-optimal performance for over-parameterized model classes interpolating the training data, as detailed in Theorem 5.1. During the rest of Section 5, we illustrated how a broad spectrum of existing regularization techniques can be shown to be proxies of this optimal regularizer. In certain restrictive, yet significant, cases, we even identified how some variations of existing regularizers can act as optimal surrogates for an interpolator’s generalization error. This analysis serves to better understand the limitations of existing regularization techniques such as parameter norm, distance from initialization or input-gradient norm.
@@
Finally, in Section 7, we established how over-parameterization is an essential requisite for achieving smooth interpolators, with smoothness being definable across various dimensions, including parameter-norm, Lipschitz constant, or the employment of invariant architectures or data augmentation. Highligting how this theoretical framework is able to establish connections among different, previously thought, unrelated approaches. We concluded the discussion by addressing the compatibility of larger model classes with smooth interpolators that exhibit superior generalization.
@@
A notable limitation of this research is its assumption of a finite model class. We believe this limitation can be overcome by employing recently introduced distribution-dependent PAC-Bayes-Chernoff bounds (Casado et al., 2024). Additionally, this study has not delved into the significant role that stochastic gradient descent (SGD) plays in identifying interpolators with minimal generalization error.
Appendix A Experimental Settings
@@
A.1 Inception architecture
The employed InceptionV3 (Szegedy et al., 2016) model used in this work is taken from Zhang et al. (2017).
A.2 Learning settings for figures

### v4

@@ 7.2 Data Augmentation
Data augmentation is a widely-used technique in machine learning that enhances the robustness and generalization ability of models, particularly in the field of computer vision (Shorten & Khoshgoftaar,, 2019). By applying various transformations like rotations, reflections, scaling, and translations to existing data samples, data augmentation artificially expands the dataset. This process helps the learned models to become invariant or less sensitive to these transformations, leading to improved performance on unseen data.
@@
Recent research has delved deeply into the theoretical foundations of data augmentation (DA), with significant contributions coming from studies such as Chen et al., (2020) and Lyle et al., (2020). These works build on the assumption than the set of transformations $G$ forms a group and that these transformations do not alter the data-generating distribution, also assuming an equality in distribution assumption. Within this context, these studies illustrate how DA can decrease variance and improve bounds on generalization error. Lyle et al., (2020) further presents a PAC-Bayes bound based on these premises, showing that data augmentation can lower the bound’s complexity measure. However, these studies encounter a critical challenge: the equality in distribution assumption does not hold in reality, as it implies that any model ${\boldsymbol{\theta}}$ would yield identical expected losses on both transformed and non-transformed input data, a premise contradicted by evidence shown in Figure 9, where two models exhibit deteriorating performance as the data is increasingly transformed. Additionally, the PAC-Bayes bound introduced by Lyle et al., (2020) is not proven to be tight for interpolators. Thus, as highlighted in the introduction, merely demonstrating that data augmentation reduces the complexity term of a bound, which might be vacuous, does not assure a lower generalization error for an interpolator.
@@
In this section, we demonstrate that our theoretical framework offers a more nuanced explanation. First, we will illustrate that data augmentation defines a new, more concentrated empirical loss. This finding is consistent with prior studies indicating that data augmentation decreases the variance of the empirical loss (Chen et al.,, 2020; Lyle et al.,, 2020). Second, we will explore how the complexity term of the PAC-Chernoff bound, introduced in Theorem 4.1, becomes smaller with the application of data augmentation. As this bound is also perfectly tight for interpolators trained on augmented datasets, we can deduce that the generalization error of interpolators under data augmentation is reduced, in opposite to the non-tight bounds used in earlier research (Lyle et al.,, 2020).
@@
As argued, for example, by Chen et al., (2020), data augmentation is equivalent to optimize, using Monte-Carlo estimates, the so-called data-augmented loss, denoted as $\ell_{G}$, which is the result of averaging the standard loss over all transformations obtained from a given input ${\boldsymbol{x}}$,
@@ 8 Over-Parameterization and Smooth Interpolation
Many theoretical contributions have established that over-parameterization is directly related to the strong performance of interpolators (Neyshabur et al.,, 2019; Bubeck et al.,, 2021). Specifically, the celebrated study by (Bubeck & Sellke,, 2023) [^3] Outstanding Paper Award at NeurIPS’21., building on an isoperimetry assumption, demonstrates that any model capable of interpolating training data beneath the noise threshold requires a (Euclidean) Lipschitz constant of the order at least $\sqrt{nd/p}$, where $d$ represents the data’s ambient dimension. We sketch this result using our notation: We say a model ${\boldsymbol{\theta}}$ is interpolating below the noise threshold if $\hat{L}(D,{\boldsymbol{\theta}})\leq L^{\star}:=\min_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}L({\boldsymbol{\theta}})$. Similarly, we denote $Lip({\boldsymbol{\theta}})$ the Lipschitz constant of the model ${\boldsymbol{\theta}}$, as defined in Equation (6). Bubeck and Sellke (2023) (Theorem 1) can be informally stated as: under an isoperimetry assumption, for any $\epsilon\in(0,L^{\star})$ and any $\delta\in(0,1)$, with high probability $1-\delta$ over $D\sim\nu^{n}$, for all ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}$,
@@
The primary advantage of the aforementioned findings are their ability to establish a link between over-parameterization, interpolation, and our prior analyses about model’s smoothness and its rate function. This enables us to show how Bubeck and Sellke (2023)’s analysis of smooth interpolation, using isoperimetry and Lipschitz continuity, represents merely one of several approximations to the characterization presented in Theorem 8.1. As described Section 6, the norm of parameters, the distance from initialization, the input-gradient norm, and the Lipschitz constant of a model, serve as proxies of the (inverse) rate function of a model and, consequently, of its smoothness. We first proceed to show how Bubeck and Sellke (2023)’s characterization of smooth interpolation can be directly deduced from Theorem 8.1 under the same isoperimetry assumption. This can also be derived under log-concave assumptions using the same techniques used in Proposition 6.5 and Equation (6).
@@
$\text{if}\quad\hat{L}(D,{\boldsymbol{\theta}})\leq\epsilon\quad\text{then}\quad Lip({\boldsymbol{\theta}})\geq\sqrt{\tfrac{nd}{2c(p\ln k-\ln\delta)}}(L^{\star}-\epsilon)\,.$
We can similarly derive a bound with a giving norm or distance-from-initialization.
@@
$\text{if}\quad\hat{L}(D,{\boldsymbol{\theta}})\leq\epsilon\quad\text{then}\quad\|{\boldsymbol{\theta}}-{\boldsymbol{\theta}}_{0}\|_{2}\geq\sqrt{\tfrac{n}{8M(p\ln k-\ln\delta)}}(L^{\star}-\epsilon)\,.$
Corollary 8.2 and 8.3 imply that, as the number of samples $n$ increases, to have smooth interpolators with a small Lipschitz constant, or a small parameter norm, or a small distance from initialization, we need to increase the number of parameters of the model class too. In any case, all these results are approximations of the general result given in Theorem 8.1. Note that a similar result could be derived in the context of input-gradient norms using Proposition 6.5.
@@
9 Related Work
In the context of uniform convergence bounds, many efforts have been made towards obtaining tighter generalization bounds (Kawaguchi et al.,, 2022; Bartlett et al.,, 2017; Neyshabur et al., 2017a, ; Golowich et al.,, 2018; Liang et al.,, 2019), by, for example, only considering the models effectively visited by the optimization algorithm. However, it is clear that these bounds are not actually meaningful (see Nagarajan & Kolter, 2019b for further references and discussions). The PAC-Bayes framework has also adapted to this new paradigm by exploiting properties of the models induced by the training data set (such as low spectral norm (Neyshabur et al., 2017b, ), noise stability (Arora et al., 2018b, ), de-randomization (Negrea et al.,, 2020), and compression (Arora et al., 2018b, ; Zhou et al.,, 2019)), but none of these bounds are shown to be empirically tight in the over-parameterized model classes and, for example, are unable to describe the interplay between invariant architectures and over-parameterization. In fact, very recently, Gastpar et al., (2024) proved that, for over-parameterized model classes, none of these bounds are tight for all data-generating distributions.
Our work mainly differs from these related works in the sense that we use oracle complexity measures, that is, PAC-Chernoff bounds and the rate function assumes access to the data-generating distribution. Consequently, we are able to tightly bound the generalization error of an interpolator, as elaborated in Section 4. This capability stems from our ability to derive robust results such as Theorem 4.6 even under unbounded losses. Furthermore, Theorem 4.6 and the PAC-Chernoff bound of Theorem 4.1 enables to leverage architectural invariances with respect to transformations inherent in the data-generating distribution.
Many other recent theoretical studies try to understand specific learning techniques in deep learning in an isolated and independent way to simplify the problem (e.g, (Vidal et al.,, 2020; Arora et al.,, 2016; Gilbert et al.,, 2017; Patel et al.,, 2016; Schoenholz et al.,, 2017; Poole et al.,, 2016; Tishby & Zaslavsky,, 2015; Arora et al., 2018a, ; Gunasekar et al.,, 2018; Hardt & Ma,, 2017; Soudry et al.,, 2018; Chen et al.,, 2020; Bubeck & Sellke,, 2023)). As a result of such isolated approach, they are not fully able, most of the times, to establish links among different learning techniques, as done in this work. Furthermore, relying on the insights provided by the field of geometric deep learning (Bronstein et al.,, 2021), our work provides a theoretical explanation, from a statistical inference point of view, on why such invariances induce smoother model classes and better generalization.
@@ 10 Conclusions and Limitations
In this work, we discussed how a growing literature (Zhang et al.,, 2017; Nagarajan & Kolter, 2019b, ; Wang et al.,, 2024; Gastpar et al.,, 2024) is suggesting that bounds that solely depend on the training data are provably vacuous for over-parameterized model classes, and points to the need of exploring novel approaches to understand generalization in deep learning. Then, we argued in favor of distribution-dependent bounds, as bounds which directly depend on the data generating distribution, instead of the training data sample. A distribution-dependent PAC-Chernoff bound, outlined in Theorem 4.1, is introduced, which shows how this bound is perfectly tight for any interpolator. Based on the complexity measure of this bound, a novel definition of smoothness is proposed in Definition 3, which is able to provide a clear answer to the previously unanswered question of which interpolators generalize more effectively.
In Section 5, we showed how PAC-Chernoff bounds are able to capture the complex phenomenon of double-descent. Our bound is able to explain why is possible to have interpolators with a smaller generalization error and, at the same time, defined by a larger number of parameters.
In Section 6, we showed how the complexity measure of our PAC-Chernoff bound defines an regularizer that provides near-optimal performance for over-parameterized model classes interpolating the training data, as detailed in Theorem 6.1. During the rest of Section 6, we illustrated how a broad spectrum of existing regularization techniques can be shown to be proxies of this optimal regularizer. In certain cases, we even identified how some variations of existing regularizers can act as optimal surrogates for an interpolator’s generalization error. This analysis serves to better understand the limitations of existing regularization techniques such as parameter norm, distance from initialization or input-gradient norm.
@@
Finally, in Section 8, we established how over-parameterization is an essential requisite for achieving smooth interpolators, with smoothness being definable across various dimensions, including parameter-norm, Lipschitz constant, or the employment of invariant architectures or data augmentation. Highlighting how this theoretical framework is able to establish connections among different, previously thought, unrelated approaches. We concluded the discussion by addressing the compatibility of larger model classes with smooth interpolators that exhibit superior generalization.
@@
10.1 Discussion of Limitations
A limitation of this research is its assumption of a finite model class. We believe this limitation can be overcome by employing recently introduced PAC-Bayes-Chernoff bounds (Casado et al.,, 2024). This PAC-Bayesian bound is also a distribution-dependent bound, relying on a similar rate function. As a result, these findings could potentially be extended to infinite model classes. However, exploring such an extension is beyond the scope of this paper. Finally, this study has not delved into the significant role that stochastic gradient descent (SGD) plays in identifying interpolators with minimal generalization error.
Another limitation of our work is the lack of explicit connections between the smoothness definition introduced in this paper and other existing definitions in the literature. For instance, smoothness is often characterized using the largest eigenvalue of the Hessian matrix (Nesterov,, 2003), the change in loss value within a perturbation radius (Keskar et al., 2017b, ), or Lipschitz continuity of the loss function (Shalev-Shwartz & Ben-David,, 2014). While our notion of smoothness is tied directly to the PAC-Chernoff framework and its distribution-dependent complexity term, future work could explore how this concept aligns or diverges from these established definitions. Establishing such connections could provide additional insights into how our framework complements or generalizes existing theoretical constructs. We have chosen to focus here on the implications of our specific definition for generalization, leaving these broader comparisons for future research. This limitation underscores the potential for follow-up studies to bridge our smoothness definition with the broader landscape of smoothness concepts in machine learning theory.
Acknowledgements
AM acknowledges funding for cloud computing from Google Cloud for Researchers program, from Grant PID2022-139293NB-C31 funded by MCIN/AEI/10.13039/501100011033 and by ERDF, a way of making Europe. LO acknowledges financial support from PID2022-139856NB-I00 funded by MCIN/ AEI / 10.13039/501100011033 / FEDER, UE and from the Autonomous Community of Madrid (ELLIS Unit Madrid).
A Experimental Settings
@@
A.1 Learning Settings for Figures
A GitHub Repository with the conducted experiments can be found in https://github.com/Ludvins/2024_PAC-Chernoff-Bound. Regarding the experimental setting of this work, we have mainly used a small InceptionV3 (Szegedy et al.,, 2016) used in Zhang et al., (2017), and Cifar10 dataset (Krizhevsky et al.,, 2009). Before showing the specific specific architecture of our small Inception, taken from Zhang et al., (2017), we need to detail some of the modules that compose it:
  1. 1.
    Convolutional module: Convolutional layer, batch-normalization and ReLU activation.
  2. 2.
    Inception module with output channels $o_{1\times 1}$ and $o_{3\times 3}$: Consists on 2 different convolutional layers, one with kernel $1\times 1$ and $o_{1\times 1}$ output channels and another with kernel $3\times 3$ and $o_{3\times 3}$ output channels. The output of this layers is then concatenated, so the total number of output channels is $o_{1\times 1}+o_{3\times 3}$.
  3. 3.
    Downsample module: Convolutional module with kernel size $3$, stride $2$ and padding 00 and MaxPooling with kernel size of $3$ and stride $2$. The outputs of these two layers is concatenated.
With these elements, the architecture of our small InceptionV3 network is
  1. 1.
    Convolutional module with $96$ output channels, kernel size $3$, stride $1$ and padding 00.
  2. 2.
    Inception Module with $o_{1\times 1}=32$ and $o_{3\times 3}=32$.
  3. 3.
    Inception Module with $o_{1\times 1}=32$ and $o_{3\times 3}=48$.
  4. 4.
    DownSample Module with $o_{3\times 3}=80$.
  5. 5.
    Inception Module with $o_{1\times 1}=112$ and $o_{3\times 3}=48$.
  6. 6.
    Inception Module with $o_{1\times 1}=96$ and $o_{3\times 3}=64$.
  7. 7.
    Inception Module with $o_{1\times 1}=80$ and $o_{3\times 3}=80$.
  8. 8.
    Inception Module with $o_{1\times 1}=48$ and $o_{3\times 3}=96$.
  9. 9.
    DownSample Module with $o_{3\times 3}=96$.
  10. 10.
    Inception Module with $o_{1\times 1}=176$ and $o_{3\times 3}=160$.
  11. 11.
    Inception Module with $o_{1\times 1}=176$ and $o_{3\times 3}=160$.
  12. 12.
    Adaptative Average Pooling layer with kernel $7\times 7$.
  13. 13.
    Fully connected layer from $16464$ to the number of classes (i.e, $10$).
Where the total number of parameters of this model is $1.814.106$.